# This work has been submitted to the IEEE Access for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. xx.xxxx/ACCESS.2024.xxxxxxx

### v2

@@ IV-B4 People Detection
A CL-based people detection framework has been developed in [350]; see Fig. 31. It employs attention maps to enable network adaptation on walking-people detection in dynamic environments, thus avoiding collision. To create attention maps from DVS data, the network is integrated with a readout mechanism. It also generalizes the unsupervised STDP from [351] to a semi-supervised learning case with teacher signals (TSs). To control overfitting, an anti-Hebbian rule (i.e., negative STDP) is applied when person is absent, while positive STDP is used when a person is detected. This work further showed the potential of implementing OCL on SNN-based systems.
@@
A summary of real-world application use-cases that will benefit from SNNs with OCL, is presented in Table VI.
TABLE VI: Summary of real-world application use-cases that will benefit from SNNs with OCL capabilities, and their current state-of-the-art works.
Applications Work CL Dataset CL Learning Neuron Hardware
(Year) Setting Approach Rule Model
Object [341] Supervised Custom Representation- Modified LIF Loihi
Recognition (2022) DVS dataset based Hebbian
Adaptive
Robots Custom
IK and PID [342] Supervised dataset from Representation- PES LIF Loihi
Control (2021) robotic arms’ based
operations
Cars [343] (2021) STBP LIF Loihi
Detection [345] (2022) Supervised N-CARS - Surrogate Gradient LIF GPU
[324] (2024) STBP LIF GPU
Road Lane [348] Supervised DET - STBP LIF Loihi
Detection (2022) [352]
Autonomous
Vehicles/
Agents SLAM [349] Unsupervised Real-time data Representation- STDP LIF n/a
(2023) from sensors based
People [350] Semi-supervised KUL- Representation- STDP LIF ReckOn
Detection (2024) UAVSAFE [353] based [354]
@@ V Open Research Challenges
Due to the huge potentials and benefits of energy-efficient SNNs with OCL capabilities, continuous advancements in the NCL field are expected in the future. Toward this, we identify several open research challenges to be addressed, as follows.
Learning Rule Developments: Leveraging SNN learning rules for enabling OCL capabilities is a major challenge. Such rules may exhibit beneficial characteristics for OCL (e.g., unsupervised local learning), but may also lack some aspects (e.g., context for quick learning convergence). Therefore, developments of appropriate learning rules and their enhancements are critical for achieving OCL capabilities. Furthermore, combining multiple learning rules may also provide promising strategies for solving multiple datasets.
Adapting the Existing CL Methods in SNN Domain: Investigating how existing CL methods in Section II-E can be adapted for SNNs is potential for combining their strengths. For instance, adapting the regularization-based methods and how it impacts the performance of SNNs trained with local learning rules, particularly in dynamic environments could be explored. The optimal memory size for replay-based method in SNNs that balances desiderata effectively, particularly in resource-constrained environments can also be explored.
@@
Balancing CL Desiderata: It is very challenging and often requires trade-offs to satisfy all CL desiderata (i.e., scalability, no/minimal usage of old data, task-agnostic, positive forward and backward transfer, controlled forgetting, and fast adaptation/recovery). For example, improving learning quality for past tasks while minimizing the use of old data is non-trivial. Therefore, it is important to prioritize the desired characteristics based on the applications and balance them during the operational lifetime within acceptable performance.
Generalization Capability: The functionality of SNN-based OCL algorithms should be applicable in broader operational settings. For instance, OCL-based SLAM systems should be able to seamlessly navigate from indoor to outdoor environments. To achieve this, multiple desiderata may need to be satisfied.
Data Acquisition Quality: In OCL scenarios, systems need to learn from unlabelled data from the environments under unsupervised settings [41, 42], which are often noisy. Therefore, input pre-processing and combining multiple high-quality event-based sensors (e.g., DVS camera) with conventional sensors (e.g., LiDAR) are beneficial to enhance the quality of input data.
@@
Hardware Deployments: Deployments of SNN-based OCL algorithms should leverage the inherent low-power and massively-parallel processing strengths of the neuromorphic hardware, thus maximizing the efficiency benefits of neuromorphic computing. For instance, algorithm mapping policies on the given hardware accelerators hold an important role in determining the processing latency and efficiency.
Evaluation and Benchmarks: Standard evaluation protocols and benchmarks are required to ensure fair and comprehensive evaluation metrics for assessing the performance of SNN-based OCL algorithms. These also include different CL scenarios and settings, as well as different metrics like accuracy, model and sample storage size, computation complexity, and power/energy efficiency.
@@ VI Conclusion
This paper provides a comprehensive review of CL foundations, analyzes state-of-the-art CL methods for both DNNs and SNNs (i.e., NCL), highlights the importance of OCL for autonomous embedded systems, then discusses how SNNs can be used for enabling OCL. Specifically, it provides a holistic view on the CL desiderata, scenarios, and settings. Then, it scrutinizes the benefits, and limitations of recent NCL advancements, to highlight the gap between the state-of-the-art and the requirements of practical applications. To address this, this paper identifies open research challenges to enable OCL using SNNs in real-world application use-cases, thereby inspiring future research towards developing energy-efficient CL systems using neuromorphic computing.

### v3

@@ IV-B4 People Detection
A CL-based people detection framework has been developed in [409]; see Fig. 37. It employs attention maps to enable network adaptation on walking-people detection in dynamic environments, thus avoiding collision. To create attention maps from DVS data, the network is integrated with a readout mechanism. It also generalizes the unsupervised STDP from [412] to a semi-supervised learning case with teacher signals (TSs). To control overfitting, an anti-Hebbian rule (i.e., negative STDP) is applied when person is absent, while positive STDP is used when a person is detected. The empirical results indicated that this SNN-STDP system achieved a peak $F_{1}$ score 19% higher than a comparable same-size CNN processing DVS frames, demonstrating its effectiveness in dynamic human detection task; see Table XIII and XIV. Future works include: a) investigating multi-object contexts with multiple walking people instead of a single person, b) implementing proposed system on neuromorphic hardware ReckOn chip [410], which supports local learning, multilayer learning, and SNN ensembles with non-spiking readouts, making it suitable for this SNN-STDP-readout approach. This work further showed the potential of implementing OCL on SNN-based systems.
@@
Figure 38: Deployment of PAIBoard in a quadruped robot for autonomous navigation, where sensor fusion (UWB + vision) and hybrid SNNs enable robust real-time tracking and obstacle avoidance [393].
IV-B5 Robotic Navigation
Recent study [393] demonstrated the use of PAIBoard neuromorphic computing platform in a quadruped robot [411] performing real-world navigation tasks such as tracking (i.e., via UWB and visual fusion) and obstacle avoidance (i.e., via depth perception) based on vision-based NN processing, as shown in Fig. 38. The system leveraged SNNs for adaptive, low-power control, underscoring the viability of hybrid SNN-ANN (HNN) platforms in autonomous sensorimotor tasks. PAIBoard achieved 90.2% accuracy on CIFAR-10 using fewer neurons and cores than TrueNorth and Tianjic, highlighting efficient resource usage. It also demonstrated $791FPS/W$ energy efficiency for image classification and consumed $12.8W$ during robot tracking and obstacle avoidance task; see Table XIII and XIV. The platform is designed to handle such complex tasks, implies a potential for OCL.
TABLE XV: Summary of open research challenges with their brief description, identified research gaps, and future directions in NCL.
Challenge Description Identified Research Gaps Future Directions
Adaptive knowledge retention and transfer Includes issues like overlapping neural representations, retaining task-specific knowledge with memory efficiency. - Insufficient exploration of NCL methods considering task-similarity for BWT.
- Limited exploration of advanced compression techniques in replay-based NCL methods.
- Developing methods with enhanced BWT leveraging similarities between tasks, while avoiding interference.
- Exploring efficient memory-augmented techniques for SNNs that preserve important synapses.
- Exploring dynamic allocation of dendritic segments only when needed to avoid overfitting.
- Designing adaptive algorithms that automatically determine the number of subspace neurons based on task complexity.
Adaptation of DNN-based CL methods in SNN domain (i.e., NCL) and Hybrid SNN-DNN Approaches The adaption of current DNN-based CL methods in SNN domain for NCL developments may lead to effective solutions.
- The integration of hybrid SNN-DNN models, leveraging the strengths of both architectures to improve CL performance.
- Limited adaptation of different CL-based methods (e.g., regularization) in SNNs, particularly for OCL scenarios.
- Limited exploration of hybrid SNN-DNN models for CL.
- Limited exploration of knowledge distillation techniques like teacher-student in SNN-based CL models.
- Exploring adaptions of CL-based methods (e.g., regularization) and their impact on SNNs trained with local learning rules in dynamic environments.
- Investigating the optimal memory size for replay-based methods in SNNs.
- Hybrid approaches where DNNs handle feature extraction while SNNs process temporal or event-driven information could enhance CL in real-world applications.
- Exploring knowledge distillation to transfer knowledge from deep networks to SNNs, enabling efficient CL while preserving performance.
Adaptation to dynamic input patterns Changes in input patterns may occur during run time, thereby requiring OCL capabilities to handle dynamic variability of input. - SNN learning quality in OCL scenarios is still a developing area. - Limited development of adaptive learning rates that are responsive to temporal correlations in spiking patterns. - Investigating SNN architectures that effectively perform weight updates to maintain performance (e.g., accuracy).
- Investigating adaptive learning and weight decay rate strategies to improve the flexibility of neuron dynamics.
- Developing methods for efficient clustering and representation learning.
- Developing adaptive algorithms to explore various configurations of dendritic segments to optimize task-specific adaptation.
Balancing CL desiderata To meet multiple, often conflicting, requirements of CL. Balancing CL desiderata remains a key optimization goal for NCL algorithms. Developing methods that balance and prioritize CL desiderata trade-offs based on the specific application needs.
Input noise robustness Real-world input data is often noisy, with more complex, non-linear patterns than AWGN. Limited exploration of diverse noise types and intensities for ensuring robustness of NCL methods against noise for OCL scenarios. - Expanding the scope of evaluation to include diverse noise types and intensities.
- Designing algorithms that automatically filter irrelevant information while emphasizing critical features, thus improving overall performance.
Data representation Most of the current NCL methods use a single encoding technique, i.e., rate coding, which neglects the temporal structure of spike trains, limiting their ability to fully exploit temporal information. - Limited exploration of NCL methods using temporal encoding with high information density.
- Lack of studies for exploring encoding methods that utilize the strengths of multiple encoding methods.
- Incorporating time-encoded signals into SNNs to enhance their learning from temporal data (e.g., TTFS encoding) to improve the network’s ability to handle time-dependent patterns.
- Developing SNNs that integrate hybrid encoding techniques for optimization of task performance, latency and energy efficiency.
Scalability Current CL methods incur significant memory and power/energy overheads when the networks get bigger, thereby posing scalability issue. - Limited studies for realizing dynamic allocation of neurons without significant computational and memory overheads.
- Lack of effective and biologically inspired synaptic reorganization (pruning and growing) mechanisms for SNNs.
- Investigating the integration of more complex neuron models (e.g., Hodgkin-Huxley model), allowing for a richer representation of neuronal activity and potentially improving the model’s predictive capabilities and scalability.
- Exploring techniques to dynamically allocate neurons for new tasks while pruning redundant ones from previous tasks.
- Investigating hierarchical or modular approaches for SNN design to enable scalable architectures for CL.
- Investigating techniques for dynamic structural plasticity, inspired by the brain’s ability to reorganize neural connections.
Learning rule developments The existing CL methods often rely on non-local learning rules and primarily use a single learning paradigm, either supervised or unsupervised, limiting performance and adaptability. - Enhancements of bio-plausible learning rules (e.g., STDP) for OCL scenarios remain under-explored.
- Limited studies in combining supervised and unsupervised settings for NCL.
Developing advanced local learning rules, such as adaptive variations of STDP, can help SNN models achieve NCL capabilities, especially in OCL scenarios.
- Developing hybrid approaches for NCL that combine different learning paradigms for training SNNs (i.e., supervised and unsupervised settings), considering OCL scenarios.
TABLE XV: Continued
Challenge Description Identified Research Gaps Future Directions
Generalization capability Current methods are limited in their adaptability to drastically changing environments or task-specific nuances. Ensuring NCL methods function effectively across diverse operational settings and tasks is a developing area. - Developing adaptive NCL methods that can dynamically adjust to diverse tasks and environmental context, such as adapting to changes from indoor to outdoor conditions.
- Expanding the evaluation of NCL methods to diverse datasets like DVS Gesture to improve generalization for complex tasks.
Evaluation datasets and benchmarks Datasets should be derived or generated to leverage unique characteristics of event-based data, which properly represent the original data. - Research on neuromorphic datasets for benchmarking NCL methods is still developing.
- Inconsistent data pre-processing (e.g., time resolution), making it difficult to compare results fairly.
- Data transformation may fail to capture temporal information, thus limiting SNN capabilities.
- Current datasets are mostly small-scale and narrowly focused, limiting the generalizability and benchmarking for SNNs.
- Developing standardized pre-processing rules to ensure fairness and consistency in dataset comparisons.
- Measuring and optimizing energy savings for showing the NC potential in practical applications.
- Constructing datasets emphasizing temporal dynamics, leveraging event-driven and spike-based data generation.
- Creating larger-scale, spatial-temporal event-based datasets for diverse tasks to enhance benchmarking and real-world applicability.
- Task-specific performance metrics are suggested, such as accuracy for classification, MAE for regression, memory ratio and power/energy overhead for scalability, and adaptability scores for functionality.
Hardware deployments - Implementing NCL methods on neuromorphic processors can maximize the efficiency benefits via massively-parallel processing and inherent low-power.
- NCL methods involve frequently updating parameters, like neuron thresholds and decay rates along with the adjustments of membrane potentials and synaptic weights.
- The transition from algorithmic level to practical implementation on hardware remains under-explored.
- Limited exploration of in-memory computing paradigms for handling frequent updates efficiently.
- Cross-integration of diverse paradigms (e.g., DNNs and SNNs) on neuromorphic chips remains under-explored.
- Developing more efficient hardware designs and leverage SNN mapping for supporting NCL methods.
- Introducing approximations in neuron parameters and synaptic weights to meet hardware constraints.
- Exploring the potential of beyond-CMOS technologies (e.g., NVM) for more energy-efficient implementation.
- Developing CIM systems with high-performance interfaces to support quick and frequent updates, minimizing latency and energy overheads.
- Developing processors for supporting the combination of unsupervised local learning with supervised global learning (e.g., combining different conventional & neuromorphic chips).
@@ V Open Research Challenges
Due to the huge potentials and benefits of energy-efficient SNNs with OCL capabilities, continuous advancements in the NCL field are expected in the future. In this section, we identify open research challenges in NCL, highlight the gaps between current research efforts, and propose future research directions; see the summary presented in Table XV. This discussion also addresses the key question Q11.
Adaptive Knowledge Retention and Transfer: Recent NCL methods mitigate CF through adaptive learning rates and weight decay coupled with STDP [42], parameter adjustments with weight potentiation and depression [43], error-corrected predictions, multi-layer architectures for maintaining distinct task representations [303], dendritic-dependent spike time delays and context-dependent gating system [356], as well as compressed LR of learned data [299]. Moreover, Hebbian learning and lateral connections, to project neuronal activity into an orthogonal subspace [365] have been utilized to preserve previous knowledge. Despite these advancements, we identify the existing research gaps and potential future directions, as follows.
  1. 1.
    Exploration of task-similarity-based BWT is limited, thus the challenge of interference in overlapping classes with high feature similarity remains open. To address this, improving task-similarity-based BWT by exploring advanced feature extraction techniques to enhance class separability and reduce interference.
  2. 2.
    In replay-based methods, increasing number of classes may result in substantial growth in storage. Meanwhile, lossy compression techniques can cause information loss that degrade the performance on previously learned task. Toward this, advanced compression techniques that preserve temporal information with efficient memory-augmentation methods may preserve synaptic connections, thus enhancing performance.
  3. 3.
    To avoid overfitting, dendritic segments can be dynamically allocated.
  4. 4.
    Finally, developing adaptive algorithms to determine subspace neuron numbers based on task complexity and optimizing computational efficiency through code enhancements or asynchronous parallel neuromorphic hardware could enable real-time adaptability for OCL.
Adaptation of Existing CL Methods in SNN Domain and Hybrid SNN-DNN approaches: Exploring how existing ANN-based CL methods can be adapted to the SNN domain, to leverage the strengths of both architectures. For instance, adapting the regularization methods and how it impacts the performance of SNNs trained with local learning rules under dynamic environments can be explored. Furthermore, the optimal memory size for replay-based method in SNNs that balances desiderata effectively under dynamic environments, can also be explored. Additionally, integration of hybrid SNN-DNN models, leveraging the strengths of both architectures, where DNNs handle feature extraction while SNNs process temporal or event-driven information to enhance CL performance in real-world applications can be explored. Moreover, knowledge distillation techniques like teacher-student to transfer knowledge from deep networks to SNNs, for efficient CL can be explored.
Adaptation to Dynamic Input Patterns: In dynamic environments, input data change unpredictably. Recent advancements leveraging STDP and active dendrites offer promising solutions. STDP adjusts synaptic strength based on spike timing, enabling efficient unsupervised learning and real-time adaptation without labeled data [42, 43]. Meanwhile, active dendrites enhance adaptability by dynamically selecting task-specific sub-networks via a gating mechanism [357, 413]. Despite these advancements, we identify that the existing techniques have not exploited temporal correlations in spiking patterns in inputs, hence hindering the network from achieving higher performance. To overcome this, there are several potential solutions, such as (1) adaptive learning with effective weight decay and weight updates, to improve the flexibility of neuron dynamics for adapting to input spikes; (2) clustering and representation learning based on the input patterns; and (3) adaptive algorithms to explore various configurations of dendritic segments in active dendrite techniques to optimize task-specific adaptation.
@@
Balancing CL Desiderata: To meet multiple, often conflicting, requirements of CL is extremely challenging, requiring trade-offs to satisfy all CL desiderata (i.e., scalability, no/minimal usage of old data, task-agnostic, positive forward and backward transfer, controlled forgetting, and fast adaptation/recovery). For example, improving learning quality for past tasks while minimizing the use of old data is non-trivial. Therefore, it is important to prioritize the desired characteristics based on the applications and balance them during the operational lifetime within acceptable performance.
Input Noise Robustness: Noise in input data, such as low contrast, background interference, or additive white Gaussian noise (AWGN), can obscure important features and significantly impair learning efficiency, particularly in dynamic environments where noise patterns vary unpredictably. A recent study addressed this issue by implementing a selective attention mechanism that focuses on task-relevant features, leading to enhanced performance in noisy conditions [102]. However, its reliance on specific noise types, such as AWGN, limits its generalizability to more complex, non-linear noise patterns. To overcome this, future research should consider various noise types and intensities, ensuring robustness in real-world scenarios. Adaptive algorithms capable of real-time adjustments to varying noise conditions may also enhance resilience. Furthermore, efficient pre-processing techniques and integration of high-resolution event-based sensors (e.g., DVS cameras) with conventional sensors (e.g., RGB camera) can also improve input data quality and robustness.
Data Representation: Current NCL methods predominantly use rate coding to represent data, which is effective in some scenarios, but overlooks the temporal structure of spike trains. This limits the advantages of temporal information. Furthermore, studies on NCL methods with temporal coding is also under-explored. To address these challenges, future research may focus on incorporating temporal coding techniques with high information density (e.g., TTFS). Additionally, advancing hybrid encoding techniques that combine multiple encoding approaches may also improve performance, while optimizing latency and energy efficiency.
Scalability: Current CL methods incur significant memory and power/energy overheads when the networks get bigger, hence posing scalability issue. However, studies for realizing dynamic allocation of neurons without significant computational and memory overheads, as well as biologically inspired synaptic reorganization mechanisms are limited. To address these challenges, several actions can be taken, as follows.
  1. 1.
    Investigating the integration of more complex neuron models (e.g., Hodgkin-Huxley model) to enable richer representation of neuronal activity, thus potentially improving performance with good scalability.
  2. 2.
    Exploring techniques to dynamically allocate neurons for new tasks while pruning redundant ones from previous tasks.
  3. 3.
    Employing hierarchical or modular SNN design to enable scalable network architectures for CL.
  4. 4.
    Investigating techniques for dynamic structural plasticity to enable reorganization of neural connections.
Learning Rule Developments: Current NCL methods often rely on non-local learning rules. While variations and enhancements of STDP for OCL scenarios have shown promise, they remain under-explored. Additionally, current approaches primarily focus on a single learning paradigm, either supervised or unsupervised, which restricts the adaptability and overall performance of NCL methods. To address these limitations, future advancements should focus on developing advanced localized learning rules, such as adaptive variations of STDP, that enable SNNs to better retain past information while achieving quick learning convergence for new data in dynamic environments. Furthermore, hybrid learning approaches that integrate supervised and unsupervised paradigms could significantly enhance the training of SNNs, improving performance and adaptability in complex, real-world scenarios.
Generalization Capability: Ensuring that NCL methods can function effectively across diverse tasks and operational settings remains an open challenge. To address this, future advancements should focus on developing adaptive NCL methods capable of dynamically adjusting to varied environmental contexts, such as transitions from indoor to outdoor conditions. For this, expanding the evaluation of NCL methods to diverse datasets is important to improve generalization for complex tasks, thereby enhancing their robustness and versatility in real-world applications.
Figure 39: Prioritized roadmap for NCL challenges in relation to real-world deployment.
Standardized Framework, Evaluation Datasets, and Benchmarks: Research on neuromorphic datasets for benchmarking NCL is still in its early stage, and the existing datasets still have limitations, such as:
  1. 1.
    Inconsistent data pre-processing practices, e.g., varying time resolutions and image compression techniques, which impede fair comparisons across methods.
  2. 2.
    Current datasets are predominantly small-scale and narrowly focused, limiting their generalizability and applicability to real-world tasks.
  3. 3.
    Data transformation from conventional to event-based data (spikes) may fail to capture temporal information, thus limiting SNN capabilities.
  4. 4.
    Most works do not specify the quantitative measurements of power/energy savings, latency and memory footprint for showing the NCL potential in practical real-world applications.
To address these limitations, following actions can be taken.
  1. 1.
    Developing standardized pre-processing protocols to ensure fairness and consistency in dataset comparisons.
  2. 2.
    Creating large-scale, spatial-temporal event-based datasets tailored for diverse tasks, apart from the classification task, would improve benchmarking and enable the broader application of NCL methods.
  3. 3.
    Constructing datasets emphasizing temporal dynamics, leveraging spike-based data generation.
  4. 4.
    Task-specific performance metrics, such as accuracy for classification, MAE for regression, memory ratio and power/energy overhead for scalability, adaptability scores for functionality and robustness under noise perturbations or quantization.
  5. 5.
    Reporting and optimizing quantitative power/energy savings, latency and memory for showing the NCL potential in practical real-world applications.
@@
Hardware Deployments: Deployments of NCL methods should leverage the inherent low-power and massively-parallel processing strengths of the neuromorphic chips, thus maximizing the efficiency benefits. For instance, algorithm mapping policies on the given hardware accelerators hold an important role in determining the processing latency and efficiency. Additionally, the frequently updating parameters such as neuron thresholds, decay rates, membrane potentials, and synaptic weights, align well with the capabilities of neuromorphic hardware. However, the hardware deployments remains under-explored. Therefore, future research should focus on the following direction.
  1. 1.
    Developing more efficient hardware designs that fully utilize SNN unique properties.
  2. 2.
    Incorporating approximations in SNN parameters and weights can also help meet hardware constraints.
  3. 3.
    Exploration of beyond-CMOS technologies (e.g., NVMs like RRAM, PCM, and MRAM) for energy-efficient hardware implementations is another promising direction.
  4. 4.
    Developing CIM systems with high-performance interfaces to support quick and frequent updates, minimizing latency and energy overheads.
  5. 5.
    Developing processors for supporting the combination of unsupervised local learning with supervised global learning (e.g., combining different conventional & neuromorphic chips).
To better align research focus with practical needs, we prioritize these challenges based on their relevance to real-world application domains such as robotics and autonomous vehicles/agents. Immediate ones include adaptive knowledge retention and transfer, input noise robustness, balancing CL desiderata and hardware-constrained deployments, each of which directly impacts system reliability and performance in dynamic environments. Medium-priority challenges such as dynamic input adaptation, scalability, hybrid SNN-DNN approaches, and the development of energy-efficient learning rules will enable broader applicability. Finally, long-term ones such as generalization capability, improved data representations, adaptation of DNN-based CL methods, and standardized benchmarking frameworks, will be essential for building robust neuromorphic intelligence. A structured roadmap prioritizing NCL challenges in relation to real-world deployment is illustrated in Fig. 39.
@@ VI Conclusion
This paper presented a comprehensive survey of CL, reviewing the state-of-the-art works in both DNN and SNN-based methods. Regarding the DNN-based methods, the survey focused on the hardware deployment challenges and discussed the need for energy-efficient CL approaches. Then, it provided an extensive technical background of low-power neuromorphic systems covering its key aspects. Regarding the SNN-based CL methods (i.e., NCL), the survey categorized them and provided their comparison with relevant DNN-based CL methods, discussed and categorized hybrid approaches and focused on efficiency enhancement techniques currently used in literature. Moreover, it included the current progress of real-world NCL applications in adaptive robots and autonomous vehicles covering a wide range of use-cases, and provided quantitative analysis. Furthermore, it reported prominent neuromorphic datasets, metrics and benchmarks, emphasizing the need for standardized benchmarks and evaluation protocols, and suggested additional metrics for NCL. Among the reviewed methods, architecture-based like DSD-SNN and SOR-SNN achieved high accuracy on image classification benchmarks like MNIST, CIFAR100 and ImageNet. Methods incorporating compressed latent replay like Replay4NCL exhibited better performance in Class-IL scenario with significant latency and energy savings in speech classification. Similarly, more biologically inspired methods such as SpikeDyn and lpSpikeCon offered significant gains in memory efficiency and inference cost for low-power embedded systems. Additionally, hybrid SNN-ANN approaches are emerging as promising directions for enabling scalable and efficient CL. Despite the notable progress in the NCL field, there is still a pressing need for ongoing and significant innovation to fully realize the potential of low-power neuromorphic systems in the design of energy-efficient CL systems. The open challenges were discussed along with future directions at the end of the survey. Notably, CL strategies developed for DNNs, such as EWC, GEM, and ER are not directly transferable to SNNs due to differences in training dynamics, data encoding, and the lack of differentiability. Addressing this gap requires the design of SNN-compatible learning rules that preserve long-term knowledge while enabling online plasticity. Moreover, effective deployment of NCL methods on neuromorphic hardware remains constrained by architectural limitations, energy budgets, and on-chip learning capabilities for OCL. Future research should prioritize hardware-aware learning algorithms and standardized benchmarks to enable fair, reproducible evaluations across both domains.
Acknowledgment
This work was partially supported by United Arab Emirates University (UPAR) with Fund code 12N169. This work was also partially supported by the NYUAD Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.